# 3D Question Answering via only 2D Vision-Language Models

### v1

@@ Abstract
Figure 1: Comparison of 3D Question Answering methods. (a): a1 for 3D-based methods; a2 and a3 for hybrid (2D+3D) methods. All of these methods require computationally intensive 3D-language alignment using point cloud data for spatial reasoning. a4 is our method that leverages pre-trained LVLMs operating solely on 2D views. The well-aligned features between 2D visual features and language in 2D LVLMs enable zero-shot 3D-QA. (b): Model comparison on the test set (with objects) of ScanQA. The upper-right corner indicates the best performance. The circle area represents the size of training data required for aligning 3D and language. The “” denotes zero-shot 3D-QA using 2D model LLAVA-OV (Li et al., 2024a). We respectively use ① uniform sampling, ② image retrieval, and ③ our cdViews, to select views as input to LLAVA-OV.
@@ 1 Introduction
The advancement of large vision-language models (LVLMs) has transformed the vision-language domain by jointly processing huge sets of vision and text training data, leading to significant breakthroughs in addressing 2D visual question answering (2D-VQA) (Shao et al., 2023; Guo et al., 2023; Lu et al., 2023). However, extending these capabilities to 3D question answering (3D-QA) has unique challenges. Unlike 2D tasks, which benefit from abundant paired training data, the 3D domain lacks large-scale datasets to learn the alignment between 3D (such as point clouds) and language (such as text descriptions of 3D scenes). Existing 3D-language models still fall short of serving as robust counterparts to the widely used 2D-language models such as LLaMA-3 (Dubey et al., 2024). Therefore, current 3D-QA methods often have to train from scratch on small-scale 3D datasets, resulting in poor model performance. In contrast, hybrid approaches leverage additional 2D information. One solution (Hong et al., 2023) is to reconstruct 3D features from the features of multiple 2D views (Figure 1 (a2)), but its performance is poor due to the technical challenge of 3D reconstruction. Another solution (Mo & Liu, 2024) is to combine 2D and 3D features as input into the model (Figure 1 (a3)). 2D features extracted from LVLMs are already well-aligned with language, but further alignment with 3D features requires careful model design and advanced training techniques. Figure 1(b) shows that hybrid methods also require extensive amounts of training data (indicated by the large circle area), which are not always available.
@@
During our preliminary trials, we identified several challenges. First, all LVLMs have a token limit, restricting the number of 2D views they can process at once. This constraint makes it crucial to carefully select the most informative views. Second, given a fixed number of views, the quality of the selected views plays a critical role. Existing methods for view selection fall into two categories: uniform sampling, which randomly selects views, and image retrieval, which selects views based on question-based retrieval (Li et al., 2022). However, both approaches have significant limitations, either being inefficient or failing to capture critical views. Specifically, as shown in Figure 2, image retrieval outperforms uniform sampling but has two major limitations. First, it prioritizes question-related views over truly essential ones for answering the question. For example, when asked “What is the black couch facing?”, the model retrieves images of the “couch” but overlooks the “coffee table”, which is the answer-related object but in the opposite view of “couch”. Second, it often selects redundant or overlapping views, causing inefficiency.
@@
We evaluate the proposed cdViews on two widely used benchmarks of 3D-QA: ScanQA (Azuma et al., 2022) and SQA (Ma et al., 2022). Our experimental results demonstrate that cdViews’s view selection significantly outperforms conventional approaches such as uniform sampling and image-text retrieval. Notably, cdViews achieves superior performance compared to models using 3D or hybrid input data. In summary, our contributions are three-fold. (1) We explore the use of 2D-only LVLM to address 3D-QA in a zero-shot manner, analyzing various view selection methods. (2) We introduce cdViews that integrates a viewSelector with a viewNMS to capture critical and diverse views. We design a viewAnnotator to generate training data for viewSelector automatically. (3) Our experiment results demonstrate that cdViews achieves state-of-the-art performance on two 3D-QA benchmarks, even surpassing the 3D or hybrid models.
@@ 2 Related Works
3D-based Methods. The 3D-based methods (Man et al., 2024a) use 3D point clouds as visual input, allowing direct processing of point cloud data to understand 3D environments. However, these methods face two challenges. First, the scarcity of 3D-language training data limits its scalability. Efforts such as 3D-VLP (Yang et al., 2024) attempt to mitigate this issue by leveraging large-scale synthetic datasets, and recent works (Zhang et al., 2024; Jin et al., 2023b; Hong et al., 2023; Zhu et al., 2023; Chen et al., 2024b) aim to unify multiple 3D tasks, such as captioning, question answering, and grounding, under a single framework. Second, using an entire 3D scene as input introduces unnecessary information for QA, distracting the model and reducing efficiency. To address this, methods such as SIG3D (Man et al., 2024a) incorporate situational awareness to focus on only relevant 3D regions guided by the language prompts (e.g., the input situation). Overall, 3D-based methods have constraints due to the lack of large-scale 3D language pretraining data. The resulting 3D-language alignment in the feature space is thus suboptimal. Besides, using entire scenes as input to answer local questions is costly and inefficient.
2D-based Methods. Recent 2D-based methods use uniformly sampled 2D views as input to 2D LVLMs (Singh et al., 2024; Zheng et al., 2024; Liu et al., 2024b), primarily focusing on evaluating the performance of 2D LVLMs on 3D-QA. They focus more on evaluating pretrained 2D LVLMs on 3D-QA tasks, rather than developing approaches to adapt and improve their performance for spatial reasoning. Some more recent works have attempted to utilize 2D views more effectively. OpenEQA (Majumdar et al., 2024), transforms visual information into textual context, such as frame-level or scene-graph captions, and then leverages LLMs to answer questions. This approach depends on whether the generated text description can accurately capture the critical visual details, which may lead to incomplete or inaccurate information.
@@
Hybrid Methods. Hybrid methods (Huang et al., 2024; Mo & Liu, 2024; Huang et al., 2023; Hong et al., 2023; Man et al., 2024b; Fu et al., 2024) leverage pre-trained 2D LVLMs to address 3D vision-language tasks in two main ways. The first approach involves mapping multi-view 2D image features (which are well-aligned with language due to 2D LVLMs) into the 3D feature space (Zhu et al., 2025; Hong et al., 2023). These mapped features can either replace original 3D features (Hong et al., 2023) or serve as complementary inputs to enhance the alignment between language and hybrid (2D+3D) features (Zhu et al., 2025). The second approach processes 2D images and 3D point clouds as parallel inputs (Mo & Liu, 2024), using complementary strengths: 2D views provide fine-grained semantic details, while 3D point clouds capture spatial awareness. Although these methods improve 3D-QA performance, they rely on explicit 3D reconstruction, needing additional models and causing more processing steps. In contrast, our method uses 2D views and feeds them into a unified 2D LVLM, which makes a simpler pipeline.
@@
Figure 4: Performance comparison of view selection methods on the validation set of ScanQA (Azuma et al., 2022). It can be observed that: 1) performance improves with an increasing number of views, peaks at a certain point, and finally declines; and 2) noticeable performance gaps arise from different view selection methods, highlighting the importance of effective view selection. An earlier peak (30.1) appears in cdViews thanks to viewNMS.
@@ 3 Preliminaries
Leveraging pre-trained 2D LVLMs in a zero-shot manner for 3D-QA tasks is promising yet underexplored. Since 2D LVLMs are fundamentally designed to process 2D images as input, we propose cdViews to efficiently select the most informative 2D views of 3D scenes. To understand the complexities in view selection, we conduct a preliminary study using intuitive view selection methods, taking LLAVA-OV (Li et al., 2024a) as the backbone and using the validation set of the ScanQA dataset (Azuma et al., 2022). Note that it requires no training data due to the zero-shot approach. In the following, we first present a problem formulation for zero-shot 3D-QA, followed by experiments using two intuitive view selection methods: uniform sampling and image retrieval.
Problem Formulation. Given a question $Q$ and a 3D scene represented by a set of 2D views =$\mathcal{M}=\{V_{1},V_{2},\dots,V_{N}\}$, each associated with a camera matrix containing the position and orientation. The view selection identifies a subset of $k$ views (that are useful to answer $Q$), denoted as $\mathcal{M}^{\prime}$:
@@
2) Image retrieval has been used in BridgeQA (Mo & Liu, 2024). Following (Mo & Liu, 2024), we use the BLIP’s image-text retrieval model (Li et al., 2022) to select views that best match the question $Q$ (option ② in Figure 3). This process can be represented as:
@@
$\hat{S}_{i}=\frac{\mathbf{{q}}\cdot\mathbf{{v}}_{i}}{||\mathbf{{q}}||||\mathbf{v}_{i}||}.$ (7)
@@
where $N^{\prime}\leq N$ is the number of views labeled as $1$ (“positive”) or 00 (“negative”).
@@
Specifically, viewNMS operates in three steps: 1) Ranking views sorts all views $\{V_{i}\}_{i=1}^{N}$ by their scores $\{\hat{S}_{i}\}_{i=1}^{N}$ in descending order, resulting in $\{V_{i_{k}}\}_{k=1}^{N}$, where $I_{i_{1}}$ is the highest-scoring view. 2) Initializing candidate views selects the highest-scoring view as the initial set =$\mathcal{M}^{\prime}=\{V_{i_{1}}\}$. 3) Adding diverse views sequentially processes the remaining views in sorted order, adding a view $V_{i_{k}}$ to the set if its distance from previously selected views exceeds a threshold $T$, formulated as:
@@
        ScanQA (Azuma et al., 2022) 3D 23.5 / 20.9 31.6 / 30.7 34.3 / 31.1 67.3 / 60.2 45.3
        SQA3D (Ma et al., 2022) 3D - - - - 47.2
        3D-LLM (Hong et al., 2023) 3D 19.1 / - 38.3 / - 35.3 / - 69.6 / - 48.1
        3D-VLP (Jin et al., 2023a) 3D 24.6 / 21.6 33.2 / 31.5 36.0 / 31.8 70.2 / 63.4 -
        3D-VisTA (Zhu et al., 2023) 3D 27.0 / 23.0 - 38.6 / 32.8 76.6 / 62.6 48.5
        SIG3D (Man et al., 2024a) 3D - - - - 52.6
        SynFormer3D (Yang et al., 2024) 3D 27.6 / 24.1 - 39.2 / 33.3 76.2 / 62.7 -
        LL3DA (Chen et al., 2024a) 3D+2D - - 38.2 / 35.2 78.2 / 70.3 -
        PQ3D (Zhu et al., 2025) 3D+2D 26.1 / 20.0 43.0 / 36.1 - 87.8 / 65.2 47.1
        BridgeQA (Mo & Liu, 2024) 3D+2D 31.3 / 30.8 34.5 / 34.4 43.3 / 41.2 83.8 / 79.3 52.9
@@
        LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ 2D 35.0 / 35.6 46.1 / 47.2 49.7 / 49.5 102.8 / 100.4 56.9
@@
Table 1: Performance comparisons with the state-of-the-art methods on the test set of ScanQA (Azuma et al., 2022) and SQA (Ma et al., 2022). For ScanQA, scores are presented in the format “with object test set” / “without object test set”. The best and second best results are in bold and underlined, and the last row shows the performance margins between LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ and the top-performing related methods.
View Distance Calculation. The core of viewNMS lies in the calculation of the view distance, i.e., $D(V_{i},V_{j})$, measuring the cameras’ position and orientation distance between $V_{i}$ and $V_{j}$. For each view, the camera parameters $[\mathbf{R}|\mathbf{t}]$ (we omit the subscript for simplicity) consist of a camera orientation $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and a camera position $\mathbf{t}\in\mathbb{R}^{3\times 1}$. The distance is calculated by combining both the orientation distance and position distance. For the orientation $\mathbf{R}$, we first convert it into a quaternion representation $\mathbf{p}=[p_{x},p_{y},p_{z},p_{w}]$ for more efficient distance calculations. Then, the orientation distance $D_{ori}(V_{i},V_{j})$ is calculated by
@@
$D_{pos}(V_{i},V_{j})=||\mathbf{t}_{i}-\mathbf{t}_{j}||,$ (11)
where |||||$\cdot$||| | ⋅ | | is the Euclidean norm.
@@ 5 Experiments
Datasets. We use ScanQA (Azuma et al., 2022) and SQA (Ma et al., 2022) in our experiments, both constructed from ScanNet dataset (Dai et al., 2017). ScanQA contains over 41K question-answer annotations across 800 indoor 3D scenes, which are divided into train, val, and test sets (with or without objects). SQA contains over 33K question-answer pairs derived from 650 indoor scenes. It encompasses a diverse range of question types, including object identification, spatial relationships, scene-level understanding, and general reasoning.
Evaluation Metrics. We adopt Exact Match (EM@1) for both datasets. EM@1 measures the proportion of cases where the top-1 predicted answers match any of the ground-truth answers. Furthermore, since the answers in ScanQA are often free-form, we use standard text similarity metrics, including BLEU-1 (Papineni et al., 2002), ROUGE-L (Lin, 2004), and CIDEr (Vedantam et al., 2015) to assess the quality of generated answers.
Implementation Details. We utilize a recent state-of-the-art LVLM, i.e., LLAVA-OV-7B (Li et al., 2024a), as the 2D LVLM for all experiments, including viewAnnotator and 3D-QA. The model remains frozen throughout all experiments. Analysis on more LVLM backbones is shown in Appendix C. The only trainable component is viewSelector, a lightweight module with a total of $5.9M$ parameters. Training of the viewSelector is conducted with a learning rate of $5\times 10^{-5}$ and a batch size of $8$. Each training iteration samples $5$ positive and $5$ negative views per instance generated by viewAnnotator. Here the number of views, e.g., $k\!=\!9$ for cdViews, is selected on the validation set (Figure 4).
Figure 6: Qualitative results for BridgeQA (Mo & Liu, 2024), LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and our final model LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$. The marks Refer to caption, Refer to caption, and Refer to caption represents the selected views respectively by three methods. We can see that cdViews captures the most critical and diverse views to answer the questions.

### v2

@@ Abstract
Figure 1: Comparison of 3D Question Answering methods. (a): a1 for 3D-based methods; a2 and a3 for hybrid (2D+3D) methods. All of these methods require computationally intensive 3D-language alignment using point cloud data for spatial reasoning. a4 is our method that leverages pre-trained LVLMs operating solely on 2D views. The well-aligned features between 2D visual features and language in 2D LVLMs enable zero-shot 3D-QA. (b): Model comparison on the test set (with objects) of ScanQA. The upper-right corner indicates the best performance. The circle area represents the size of training data required for aligning 3D and language. The “” denotes zero-shot 3D-QA using 2D model LLAVA-OV (li2024llava). We respectively use ① uniform sampling, ② image retrieval, and ③ our cdViews, to select views as input to LLAVA-OV.
@@ 1 Introduction
The advancement of large vision-language models (LVLMs) has transformed the vision-language domain by jointly processing huge sets of vision and text training data, leading to significant breakthroughs in addressing 2D visual question answering (2D-VQA) (shao2023prompting; guo2023images; lu2023multi). However, extending these capabilities to 3D question answering (3D-QA) has unique challenges. Unlike 2D tasks, which benefit from abundant paired training data, the 3D domain lacks large-scale datasets to learn the alignment between 3D (such as point clouds) and language (such as text descriptions of 3D scenes). Existing 3D-language models still fall short of serving as robust counterparts to the widely used 2D-language models such as LLaMA-3 (dubey2024llama). Therefore, current 3D-QA methods often have to train from scratch on small-scale 3D datasets, resulting in poor model performance. In contrast, hybrid approaches leverage additional 2D information. One solution (hong20233d) is to reconstruct 3D features from the features of multiple 2D views (Figure 1 (a2)), but its performance is poor due to the technical challenge of 3D reconstruction. Another solution (mo2024bridging) is to combine 2D and 3D features as input into the model (Figure 1 (a3)). 2D features extracted from LVLMs are already well-aligned with language, but further alignment with 3D features requires careful model design and advanced training techniques. Figure 1(b) shows that hybrid methods also require extensive amounts of training data (indicated by the large circle area), which are not always available.
@@
During our preliminary trials, we identified several challenges. First, all LVLMs have a token limit, restricting the number of 2D views they can process at once. This constraint makes it crucial to carefully select the most informative views. Second, given a fixed number of views, the quality of the selected views plays a critical role. Existing methods for view selection fall into two categories: uniform sampling, which randomly selects views, and image retrieval, which selects views based on question-based retrieval (li2022blip). However, both approaches have significant limitations, either being inefficient or failing to capture critical views. Specifically, as shown in Figure 2, image retrieval outperforms uniform sampling but has two major limitations. First, it prioritizes question-related views over truly essential ones for answering the question. For example, when asked “What is the black couch facing?”, the model retrieves images of the “couch” but overlooks the “coffee table”, which is the answer-related object but in the opposite view of “couch”. Second, it often selects redundant or overlapping views, causing inefficiency.
@@
We evaluate the proposed cdViews on two widely used benchmarks of 3D-QA: ScanQA (azuma2022scanqa) and SQA (ma2022sqa3d). Our experimental results demonstrate that cdViews’s view selection significantly outperforms conventional approaches such as uniform sampling and image-text retrieval. Notably, cdViews achieves superior performance compared to models using 3D or hybrid input data. In summary, our contributions are three-fold. (1) We explore the use of 2D-only LVLM to address 3D-QA in a zero-shot manner, analyzing various view selection methods. (2) We introduce cdViews that integrates a viewSelector with a viewNMS to capture critical and diverse views. We design a viewAnnotator to generate training data for viewSelector automatically. (3) Our experiment results demonstrate that cdViews achieves state-of-the-art performance on two 3D-QA benchmarks, even surpassing the 3D or hybrid models.
@@ 2 Related Works
3D-based Methods. The 3D-based methods (man2024situational) use 3D point clouds as visual input, allowing direct processing of point cloud data to understand 3D environments. However, these methods face two challenges. First, the scarcity of 3D-language training data limits its scalability. Efforts such as 3D-VLP (yang20243d) attempt to mitigate this issue by leveraging large-scale synthetic datasets, and recent works (zhang2024vision; jin2023context; hong20233d; zhu20233dvista; chen2024grounded) aim to unify multiple 3D tasks, such as captioning, question answering, and grounding, under a single framework. Second, using an entire 3D scene as input introduces unnecessary information for QA, distracting the model and reducing efficiency. To address this, methods such as SIG3D (man2024situational) incorporate situational awareness to focus on only relevant 3D regions guided by the language prompts (e.g., the input situation). Overall, 3D-based methods have constraints due to the lack of large-scale 3D language pretraining data. The resulting 3D-language alignment in the feature space is thus suboptimal. Besides, using entire scenes as input to answer local questions is costly and inefficient.
2D-based Methods. Recent 2D-based methods use uniformly sampled 2D views as input to 2D LVLMs (singh2024evaluating; zheng2024towards; liu2024llavanext), primarily focusing on evaluating the performance of 2D LVLMs on 3D-QA. They focus more on evaluating pretrained 2D LVLMs on 3D-QA tasks, rather than developing approaches to adapt and improve their performance for spatial reasoning. Some more recent works have attempted to utilize 2D views more effectively. OpenEQA (majumdar2024openeqa), transforms visual information into textual context, such as frame-level or scene-graph captions, and then leverages LLMs to answer questions. This approach depends on whether the generated text description can accurately capture the critical visual details, which may lead to incomplete or inaccurate information.
@@
Hybrid Methods. Hybrid methods (huang2024chat; mo2024bridging; huang2023embodied; hong20233d; man2024lexicon3d; fu2024scene) leverage pre-trained 2D LVLMs to address 3D vision-language tasks in two main ways. The first approach involves mapping multi-view 2D image features (which are well-aligned with language due to 2D LVLMs) into the 3D feature space (zhu2025unifying; hong20233d). These mapped features can either replace original 3D features (hong20233d) or serve as complementary inputs to enhance the alignment between language and hybrid (2D+3D) features (zhu2025unifying). The second approach processes 2D images and 3D point clouds as parallel inputs (mo2024bridging), using complementary strengths: 2D views provide fine-grained semantic details, while 3D point clouds capture spatial awareness. Although these methods improve 3D-QA performance, they rely on explicit 3D reconstruction, needing additional models and causing more processing steps. In contrast, our method uses 2D views and feeds them into a unified 2D LVLM, which makes a simpler pipeline.
@@
Figure 4: Performance comparison of view selection methods on the validation set of ScanQA (azuma2022scanqa). It can be observed that: 1) performance improves with an increasing number of views, peaks at a certain point, and finally declines; and 2) noticeable performance gaps arise from different view selection methods, highlighting the importance of effective view selection. An earlier peak (30.1) appears in cdViews thanks to viewNMS.
@@ 3 Preliminaries
Leveraging pre-trained 2D LVLMs in a zero-shot manner for 3D-QA tasks is promising yet underexplored. Since 2D LVLMs are fundamentally designed to process 2D images as input, we propose cdViews to efficiently select the most informative 2D views of 3D scenes. To understand the complexities in view selection, we conduct a preliminary study using intuitive view selection methods, taking LLAVA-OV (li2024llava) as the backbone and using the validation set of the ScanQA dataset (azuma2022scanqa). Note that it requires no training data due to the zero-shot approach. In the following, we first present a problem formulation for zero-shot 3D-QA, followed by experiments using two intuitive view selection methods: uniform sampling and image retrieval.
Problem Formulation. Given a question $Q$ and a 3D scene represented by a set of 2D views $\mathcal{M}=\{V_{1},V_{2},\dots,V_{N}\}$, each associated with a camera matrix containing the position and orientation. The view selection identifies a subset of $k$ views (that are useful to answer $Q$), denoted as $\mathcal{M}^{\prime}$:
@@
2) Image retrieval has been used in BridgeQA (mo2024bridging). Following (mo2024bridging), we use the BLIP’s image-text retrieval model (li2022blip) to select views that best match the question $Q$ (option ② in Figure 3). This process can be represented as:
@@
[…]
@@
where $N^{\prime}\leq N$ is the number of views labeled as $1$ (“positive”) or 0 (“negative”).
@@
Specifically, viewNMS operates in three steps: 1) Ranking views sorts all views $\{V_{i}\}_{i=1}^{N}$ by their scores $\{\hat{S}_{i}\}_{i=1}^{N}$ in descending order, resulting in $\{V_{i_{k}}\}_{k=1}^{N}$, where $I_{i_{1}}$ is the highest-scoring view. 2) Initializing candidate views selects the highest-scoring view as the initial set $\mathcal{M}^{\prime}=\{V_{i_{1}}\}$. 3) Adding diverse views sequentially processes the remaining views in sorted order, adding a view $V_{i_{k}}$ to the set if its distance from previously selected views exceeds a threshold $T$, formulated as:
@@
     ScanQA (azuma2022scanqa) 3D 23.5 / 20.9 31.6 / 30.7 34.3 / 31.1 67.3 / 60.2 45.3
     SQA3D (ma2022sqa3d) 3D - - - - 47.2
     3D-LLM (hong20233d) 3D 19.1 / - 38.3 / - 35.3 / - 69.6 / - 48.1
     3D-VLP (Jin_2023_CVPR) 3D 24.6 / 21.6 33.2 / 31.5 36.0 / 31.8 70.2 / 63.4 -
     3D-VisTA (zhu20233dvista) 3D 27.0 / 23.0 - 38.6 / 32.8 76.6 / 62.6 48.5
     SIG3D (man2024situational) 3D - - - - 52.6
     SynFormer3D (yang20243d) 3D 27.6 / 24.1 - 39.2 / 33.3 76.2 / 62.7 -
     LL3DA (chen2024ll3da) 3D+2D - - 38.2 / 35.2 78.2 / 70.3 -
     PQ3D (zhu2025unifying) 3D+2D 26.1 / 20.0 43.0 / 36.1 - 87.8 / 65.2 47.1
     BridgeQA (mo2024bridging) 3D+2D 31.3 / 30.8 34.5 / 34.4 43.3 / 41.2 83.8 / 79.3 52.9
@@
     LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ 2D 35.0 / 35.6 46.1 / 47.2 49.7 / 49.5 102.8 / 100.4 56.8
@@
Table 1: Performance comparisons with the state-of-the-art methods on the test set of ScanQA (azuma2022scanqa) and SQA (ma2022sqa3d). For ScanQA, scores are presented in the format “with object test set” / “without object test set”. The best and second best results are in bold and underlined, and the last row shows the performance margins between LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ and the top-performing related methods.
View Distance Calculation. The core of viewNMS lies in the calculation of the view distance, i.e., $D(V_{i},V_{j})$, measuring the cameras’ position and orientation distance between $V_{i}$ and $V_{j}$. For each view, the camera parameters $\mathbf{R}|\mathbf{t}]$ (we omit the subscript for simplicity) consist of a camera orientation $\mathbf{R}\in\mathbb{R}^{3\times 3}$ and a camera position $\mathbf{t}\in\mathbb{R}^{3\times 1}$. The distance is calculated by combining both the orientation distance and position distance. For the orientation $\mathbf{R}$, we first convert it into a quaternion representation $\mathbf{p}=[p_{x},p_{y},p_{z},p_{w}]$ for more efficient distance calculations. Then, the orientation distance $D_{ori}(V_{i},V_{j})$ is calculated by
@@
[…]
where $||\cdot||$ is the Euclidean norm.
@@ 5 Experiments
Datasets. We use ScanQA (azuma2022scanqa) and SQA (ma2022sqa3d) in our experiments, both constructed from ScanNet dataset (dai2017scannet). ScanQA contains over 41K question-answer annotations across 800 indoor 3D scenes, which are divided into train, val, and test sets (with or without objects). SQA contains over 33K question-answer pairs derived from 650 indoor scenes. It encompasses a diverse range of question types, including object identification, spatial relationships, scene-level understanding, and general reasoning.
Evaluation Metrics. We adopt Exact Match (EM@1) for both datasets. EM@1 measures the proportion of cases where the top-1 predicted answers match any of the ground-truth answers. Furthermore, since the answers in ScanQA are often free-form, we use standard text similarity metrics, including BLEU-1 (papineni2002bleu), ROUGE-L (lin2004rouge), and CIDEr (vedantam2015cider) to assess the quality of generated answers.
Implementation Details. We utilize a recent state-of-the-art LVLM, i.e., LLAVA-OV-7B (li2024llava), as the 2D LVLM for all experiments, including viewAnnotator and 3D-QA. The model remains frozen throughout all experiments. Analysis on more LVLM backbones is shown in Appendix C. The only trainable component is viewSelector, a lightweight module with a total of $5.9M$ parameters. Training of the viewSelector is conducted with a learning rate of $5\times 10^{-5}$ and a batch size of $8$. Each training iteration samples $5$ positive and $5$ negative views per instance generated by viewAnnotator. Here the number of views, e.g., k=9k\!=\!9 for cdViews, is selected on the validation set (Figure 4).
Figure 6: Qualitative results for BridgeQA (mo2024bridging), LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and our final model LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$. The marks Refer to caption, Refer to caption, and Refer to caption represents the selected views respectively by three methods. We can see that cdViews captures the most critical and diverse views to answer the questions.